OpenStudy (anonymous):

a mule and donkey were going to the market laden with wheat the mule said"if you give me one measure, i should carry twice as much as you, but if i give you one, we should have equal burdens." what were there burdens?

OpenStudy (mathstudent55):

The mule has m amount, and the donkey has d amount before they move the amounts around. mule ---> m donkey ---> d If the donkey gives one measure, then the donkey will have d - 1 amount and the mule will have m + 1 amount. Now they have: mule ---> m + 1 donkey ---> d - 1 We are told this means the mule has twice the donkey's amount, so m + 1 = 2(d - 1) Equation 1

OpenStudy (mathstudent55):

We now have one equation is two variables. We need a second equation, and we are given another piece of information that will let us find the second equation.

OpenStudy (anonymous):

pls give answer

OpenStudy (mathstudent55):

If the mule gives one measure to the donkey, the burdens are equal. mule ---> m - 1 donkey ---> d + 1 m - 1 = d + 1 Equation 2

OpenStudy (mathstudent55):

Now we solve the two equations in a system of equations. m + 1 = 2(d - 1) ---> m + 1 = 2d - 2 ---> m = 2d - 3 m - 1 = d + 1 m - 1 = d + 1 m = d + 2 Since both equations are solved for m, we set them equal to each other. 2d - 3 = d + 2 d = 5 m - 1 = d + 1 m - 1 = 5 + 1 m - 1 = 6 m = 7 The mule has 7 measures and the donkey has 5 measures.

OpenStudy (anonymous):

yessssss